OpenStudy (nali):

plz help... use the point on the line and the slope of the line to find the general form of the equation of the line (0, 7/8) m=-4/5

OpenStudy (nali):

general form ax+by+c = 0

OpenStudy (turingtest):

start from point-slope form given the point \((x_0,y_0)\) and the slope \(m\) we get the formula\[y-y_0=m (x-x_0)\]distribute the given \(m\) and rearrange the expression so that it is equal to zero, giving the equation in general form.

OpenStudy (nali):

i know thsi part but i do not know how to change it to general form afterwards

OpenStudy (turingtest):

I can always rearrange this expression to solve for for 0;\[y-y_0=m(x-x_0)\]\[y-y_0=mx-mx_0\]\[y-mx+mx_0-y_0=0\]notice that the last two terms are a constant, which will be c in your equation\[y-mx+(mx_0-y_0)=0\]which is the same as\[ay+bx+c=0\]

OpenStudy (nali):

can u plz use the numbers instead of the variables so i could understand

OpenStudy (turingtest):

if the point were\[(x_0,y_0)=(1,3)\]and the slope were \(m=2\) it would be\[y-y_0=m(x-x_0)\]\[y-3=2(x-1)\]\[y-3=2x-2\]\[y+2x-1=0\]

OpenStudy (nali):

is there a rule for having the general form other than switching number and variables ?around

OpenStudy (nali):

numbers*

OpenStudy (turingtest):

the rule is that the general form is the form you wrote: ax+by+c=0 if the equation looks like that, it's in general form; that's the definition

OpenStudy (turingtest):

you also want to eliminate any fractions; so you may need to multiply by something you may have to do slightly different algebra tricks to get it from one form to another in this case you are given a point and a slope, so you have to use that info if you know point-slope form well enough you can save yourself some time, but you are safer starting from what I suggested I think....

OpenStudy (nali):

that was my main concern ^^ so i got to this point and i solved for it y+ 4/5x -7/8 = 0 now how do i do the cancellation of the fraction part?

OpenStudy (apoorvk):

for the fraction part okay. just multiply both sides of the whole equation by 5 first, and see what happens.

OpenStudy (nali):

ok i got 5y+ 4x -35/8 = 0

OpenStudy (apoorvk):

great now do the same, this time multiply every thing by 8.

OpenStudy (saifoo.khan):

@Nali where are you stuck now?

OpenStudy (nali):

i was stuck in the cancellation part that will give e big numbers 40y + 32x - 35= 0

OpenStudy (nali):

it is ok thanks to all